# Accelerating PDE Data Generation via Differential Operator Action in Solution Space

### v1

@@ 1 Introduction
Figure 1: Above. Generation process of the PDE dataset. 1. Produce collection of random parameters derived from PDE. 2. Generate the relevant PDE using these parameters 3. Convert the PDE into linear equation systems using discretization methods. 4. Solve linear equations based on input parameters 5. Acquire solutions for the linear systems and translate them into solutions for the PDEs. 6. Compile the data into a dataset. Below. The generation cost of DiffOAS and GMRES varies with the dimension of the linear system. The GMRES allows for different error tolerances, while our DiffOAS keeps the machine precision of $10^{-16}$. It can be observed that DiffOAS significantly speeds up the data generation process, achieving a speedup of up to $70,000$ times.
@@
Figure 1: Above. Generation process of the PDE dataset. 1. Produce collection of random parameters derived from PDE. 2. Generate the relevant PDE using these parameters 3. Convert the PDE into linear equation systems using discretization methods. 4. Solve linear equations based on input parameters 5. Acquire solutions for the linear systems and translate them into solutions for the PDEs. 6. Compile the data into a dataset. Below. The generation cost of DiffOAS and GMRES varies with the dimension of the linear system. The GMRES allows for different error tolerances, while our DiffOAS keeps the machine precision of $10^{-16}$. It can be observed that DiffOAS significantly speeds up the data generation process, achieving a speedup of up to $70,000$ times.
@@
However, the long running time and computation costs of generating the training datasets pose a great challenge to the training of NOs. Firstly, real-world applications often involve various types of PDEs, and the training of NO for a certain type of PDEs requires a large number of training instances. For example, when training a Fourier Neural Operator (FNO) (Li et al., 2020) for Darcy flow problems, it is common to require thousands of PDEs and their corresponding solutions under diverse initial conditions. Secondly, acquiring the corresponding solution functions of PDEs as labels for training the NO network poses a further obstacle. Existing algorithms typically rely on traditional PDE solvers, such as the finite difference method (FDM) (Strikwerda, 2004) and finite element method (FEM) (LeVeque, 2002). These algorithms, demonstrated in Figure LABEL:flow_chart, involve solving large-scale linear equation systems in the $4$-th step ’Solver’, with a high computational complexity of $O(n^{3})$, where $n$ represents the dimension of the linear system. Consequently, utilizing such algorithms can be time-consuming, with the computational cost of the ’Solver’ module often accounting for over 95% of the entire data generation process (Hughes, 2012). Thirdly, solving large-scale linear systems often involves the use of iterative methods. However, due to the presence of termination conditions, these methods inevitably introduce errors, which can potentially result in a degradation of the performance of the NO network. Furthermore, as shown in Figure 1, increasing the accuracy of solving linear systems leads to increased time costs. Therefore, these challenges of data generation have significantly hindered the real-world applications of NOs (Zhang et al., 2023).
To address the aforementioned challenges, we propose a novel and efficient PDE data generation method, named Differential Operator Action in Solution space (DiffOAS). DiffOAS has two key advantages: it accelerates the data generation process and enhances the precision of the generated data simultaneously. DiffOAS replaces the process of solving linear systems in the ’Solver’ module in Figure LABEL:flow_chart with operator actions. Initially, we generate a set of PDE solution functions that comply with the actual physical contexts, which serve as basis functions for the solution space. These basis functions are then appropriately combined to satisfy the PDE conditions and generate solution functions. The discretized operator and solution functions are used as inputs in the ’Solver’, where the corresponding differential operators are applied to calculate the remaining data that satisfies PDE constraints. DiffOAS utilizes operator actions to avoid the process of solving linear equation systems, reducing the computational complexity by one order. As shown in Figure 1, the DiffOAS method can significantly accelerate PDE data generation, achieving a speedup of up to 300 times.
@@ 4 Method
In existing algorithms, the dataset generation process began with the random generation of other parameters for PDEs, followed by employing traditional PDE solvers to determine the corresponding solution functions $u(\bm{x})$. This approach necessitated the ’Solver’ module in Figure LABEL:flow_chart to solve large linear systems of equations involving substantial matrices, as mentioned in the Preliminaries 3.1. As mentioned in Figure 1, existing method requires a substantial amount of time and often becomes a bottleneck in the dataset generation process.
@@ 4.1 Solution Functions Generation
DiffOAS method specifically generates a series of basis functions either through a designated distribution that satisfies actual physical contexts (typically choosing between 10 to 50, taking $N_{basis}$ as an example). These basis functions form the foundational elements in the solution space for that particular physical distribution. We randomly weight and normalize these basis functions using Gaussian distribution. Additionally, we introduce a noise element $\epsilon$ that maintains the boundary conditions unaltered (further details are provided in the Appendix B.2), resulting in a novel solution function $u_{new}(\bm{x})$.
@@ B.2.3 Solute Diffusion in Porous Media
[…]

### v2

@@ 1 Introduction
Figure 1: Above. Generation process of the PDE dataset. 1. Produce collection of random parameters derived from PDE. 2. Generate the relevant PDE using these parameters 3. Convert the PDE into linear equation systems using discretization methods. 4. Solve linear equations based on input parameters 5. Acquire solutions for the linear systems and translate them into solutions for the PDEs. 6. Compile the data into a dataset. Below. The generation cost of DiffOAS and GMRES varies with the dimension of the linear system. GMRES curves in the graph represent different truncation errors, where GMRES $10^{-5}$ indicates the algorithm’s time cost with a truncation error of $10^{-5}$. In contrast, our DiffOAS maintains machine precision of $10^{-16}$. It can be observed that DiffOAS significantly speeds up the data generation process, achieving a speedup of up to $70,000$ times.
@@
However, the long running time and computation costs of generating the training datasets pose a great challenge to the training of NOs. Firstly, real-world applications often involve various types of PDEs, and the training of NO for a certain type of PDEs requires a large number of training instances. For example, when training a Fourier Neural Operator (FNO) (Li et al., 2020) for Darcy flow problems, it is common to require thousands of PDEs and their corresponding solutions under diverse initial conditions. Secondly, acquiring the corresponding solution functions of PDEs as labels for training the NO network poses a further obstacle. Existing algorithms typically rely on traditional PDE solvers, such as the finite difference method (FDM) (Strikwerda, 2004) and finite element method (FEM) (LeVeque, 2002). These algorithms, demonstrated in Figure 1, involve solving large-scale linear equation systems in the $4$-th step Solver, with a high computational complexity of $O(n^{3})$, where $n$ represents the dimension of the linear system. Consequently, utilizing such algorithms can be time-consuming, with the computational cost of the Solver module often accounting for over 95% of the entire data generation process (Hughes, 2012). Thirdly, solving large-scale linear systems often involves the use of iterative methods. However, due to the presence of termination conditions, these methods inevitably introduce errors, which can potentially result in a degradation of the performance of the NO network. Furthermore, as shown in Figure 1, increasing the accuracy of solving linear systems leads to increased time costs. Therefore, these challenges of data generation have significantly hindered the real-world applications of NOs (Zhang et al., 2023).
To address the aforementioned challenges, we propose a novel and efficient PDE data generation method, named Differential Operator Action in Solution space (DiffOAS). DiffOAS has two key advantages: it accelerates the data generation process and enhances the precision of the generated data simultaneously. DiffOAS replaces the process of solving linear systems in the Solver module in Figure 1 with operator actions. Initially, we generate a set of PDE solution functions that comply with the actual physical contexts, which serve as basis functions for the solution space. These basis functions are then appropriately combined to satisfy the PDE conditions and generate solution functions. The discretized operator and solution functions are used as inputs in the Solver, where the corresponding differential operators are applied to calculate the remaining data that satisfies PDE constraints. DiffOAS utilizes operator actions to avoid the process of solving linear equation systems, reducing the computational complexity by one order. As shown in Figure 1, the DiffOAS method can significantly accelerate PDE data generation, achieving a speedup of up to 300 times.
@@ 4 Method
In existing algorithms, the dataset generation process began with the random generation of other parameters for PDEs, followed by employing traditional PDE solvers to determine the corresponding solution functions $u(\bm{x})$. This approach necessitated the Solver module in Figure 1 to solve large linear systems of equations involving substantial matrices, as mentioned in the Preliminaries 3.1. As mentioned in Figure 1, existing method requires a substantial amount of time and often becomes a bottleneck in the dataset generation process.
@@ 4.1 Solution Functions Generation
DiffOAS method specifically generates a series of basis functions either through a designated distribution that satisfies actual physical contexts (typically choosing between 10 to 50, taking $N_{basis}$ as an example). This paper utilizes existing method to generate these basis functions. These basis functions form the foundational elements in the solution space for that particular physical distribution. We randomly weight and normalize these basis functions using Gaussian distribution. Additionally, we introduce a noise element $\epsilon$ that maintains the boundary conditions unaltered (further details are provided in the Appendix B.2), resulting in a novel solution function $u_{new}(\bm{x})$.
@@ 5.2 Error Analysis
To demonstrate how larger data errors can impact model training, we conducted additional experiments detailed in Appendix C.2, where we trained models using datasets with varying error levels.
@@ 6.3 Data Generation Time And Accuracy
Our method tackles a common scenario: solving systems of linear equations involving large, sparse, non-symmetric matrices. The GMRES algorithm is the mainstream approach for addressing such problems. However, for certain types of equations, specifically when the system is symmetric positive definite, numerical algorithms like MINRES and Conjugate Gradient should also be considered for baseline comparisons. Consequently, we have included comparative experiments with these baselines in Appendix C.1.
@@ B.2.3 Solute Diffusion in Porous Media
This equation is key for modeling solute movement in heterogeneous porous media. For our experimental setup, the diffusion coefficient $k(x,y)$ is generated by $10*GRF(\tau=3,\alpha=2)$, $f(x,y)$ is generated by $GRF(\tau=3,\alpha=2)$ and the influence of internal or external sources/sinks $q(x,y)$ is generated by uniform distribution $U[0,1]$.
@@ B.4 Ablation
Appendix C Additional Experiments
C.1 Time comparison Experiments
We conducted comparative experiments involving the SKR algorithm, MINRES, and Conjugate Gradient algorithms (Wang et al., 2024; Paige & Saunders, 1975; Hestenes et al., 1952). MINRES and Conjugate Gradient algorithms are often applied to specific matrix types; for instance, the Conjugate Gradient method is tailored for symmetric positive-definite matrices. To maintain consistency in presenting results, we focused on scenarios involving symmetric positive-definite matrices:
Table 5: Considering the Darcy flow problem, appropriate boundary conditions are selected to ensure that the corresponding linear equation system is symmetric positive-definite. Each corresponding matrix has a size of 62500, and the table records the time taken by different algorithms to generate 1000 training data sets. The DiffOAS method’s time denotes the duration for generating training data after obtaining basis functions.
      Method       Time(s)
      DiffOAS(scipy)       3.893E-1
      SKR(C)       1.283E3
      GMRES(scipy)       1.974E4
      MINRES(scipy)       1.149E2
      Conjugate Gradient(scipy)       1.426E3
The experimental results show that the DiffOAS algorithm significantly outperforms other numerical algorithms in terms of computational time required for data generation. MINRES and Conjugate Gradient, optimized for specific matrix types, are faster than GMRES. Due to the large condition number of the matrices corresponding to the equations, MINRES demonstrates faster speeds than Conjugate Gradient. The SKR algorithm surpasses GMRES by optimizing the solution for multiple related linear equation systems. However, since it’s not optimized for the symmetry of matrices, its speed still lags behind MINRES. Even though our method is not tailored for specific matrix types, it still generates data faster than numerically optimized algorithms (MINRES and Conjugate Gradient) designed for specific matrix types.
It is worth noting that our method can also be combined with the SKR algorithm or other numerical algorithms. Our method mainly consists of two parts: basis function generation and training data generation after obtaining the basis functions. Through the following experiments, it can be observed that the time cost of generating data with the DiffOAS algorithm after obtaining basis functions is much lower than that of all existing methods, including the SKR algorithm. Therefore, the basis function generation becomes the primary time cost of DiffOAS. The SKR algorithm significantly improves the efficiency of generating basis functions. Due to the SKR algorithm’s ability to significantly enhance the efficiency of basis function generation, its integration also leads to notable performance improvement in our method.
C.2 Error Experiments
We evaluated models trained with datasets of varying accuracies generated by the GMRES algorithm to demonstrate the impact of low-precision data on model training.
The experimental results indicate that when there are significant data accuracy errors, the errors noticeably impact the performance of the model. This suggests that obtaining high-accuracy data has a significant impact on the quality of the dataset.
Table 6: For the Darcy flow problem with a matrix dimension of 10,000, FNO was used as the testing model. The training outcomes of models trained on datasets generated by the GMRES algorithm, featuring truncation errors of $10^{-1}$, $10^{-2}$, and $10^{-5}$, were evaluated.
      Train number       tol=1E-1       tol=1E-2       tol=1E-5
      100       1.25E-1       4.54E-2       4.41E-2
      500       1.20E-1       1.14E-2       6.87E-3
      1000       1.20E-1       1.04E-2       4.56E-3
      5000       1.20E-1       9.60E-3       1.69E-3
      10000       1.20E-1       9.43E-3       1.15E-3
      20000       1.20E-1       9.85E-3       1.01E-3